# Improving governance outcomes through AI documentation: Bridging theory and practice

### v1

@@ Abstract.
Documentation plays a crucial role in both external accountability and internal governance of AI systems. Although there are many proposals for documenting AI data, models, systems, and methods, the ways these practices enhance governance as well as the challenges practitioners and organizations face with documentation remain underexplored. In this paper, we analyze 37 proposed documentation frameworks and 21 empirical studies evaluating their use. We identify potential hypotheses about how documentation can strengthen governance—such as informing stakeholders about AI risks and usage, fostering collaboration, encouraging ethical reflection, and reinforcing best practices. However, empirical evidence shows that practitioners often encounter obstacles that prevent documentation from achieving these goals. We also highlight key considerations for organizations when designing documentation, such as determining the appropriate level of detail and balancing automation in the process. Finally, we offer recommendations for further research and for implementing effective documentation practices in real-world contexts.
@@ 1. Introduction
Transparency in AI is widely regarded as essential for calibrating trust and supporting accountability (Liao and Vaughan, 2023). By providing stakeholders with a clear view of an AI system’s composition, operation, and development process, transparency allows for informed oversight and critical evaluation. When AI systems are thoroughly documented, these documentation artifacts offer invaluable insights into important system features such as training data, algorithms, and risk management strategies. This information can support two broad groups of external stakeholders. Documentation can help downstream deployers and users understand how the system functions and what risks it conveys. It can also assist policymakers and researchers in holding companies accountable for the negative consequences of their AI technologies.
Yet the importance of documentation extends beyond external accountability.111Though the concepts of documentation and transparency are often blurred, complete documentation about an AI system does not guarantee that information will be made available to interested stakeholders, nor do transparency artifacts always reflect the full or most relevant details of an AI system even though such details may have been captured in internal documents. We distinguish between the concepts in this report in order to provide greater clarity about the presumptive goals of documentation and an exploration of what practices may more likely support those objectives Internally, documentation serves as a critical tool for managing AI systems throughout their lifecycle for a wide range of stakeholders. For example, documentation could help downstream technical practitioners understand the strengths, limitations, and risks of training a model on a given dataset. It might also help compliance and governance teams assess a potential system use case’s compliance with company policies or legal requirements. Or the process of producing documentation might lead practitioners to make different decisions about how to sufficiently mitigate potential harms. As such, documentation is a foundational building block of effective AI governance.
@@
In this paper, we explore how documentation processes and artifacts can best support AI governance and risk management goals. Our research draws from a qualitative analysis of academic and gray literature,222Gray literature refers to publications that have not been peer-reviewed, but that present detailed theories or research findings including 37 proposed documentation frameworks and 21 papers with relevant empirical findings. From this analysis, we identify four mechanisms through which documentation is hypothesized to improve AI governance:
Informing stakeholders about responsible use,
Facilitating collaboration on AI risks,
Encouraging ethical deliberation, and
Improving overall governance and development practices.
We also examine the practical barriers that may hinder the effectiveness of these mechanisms and the considerations or trade-offs organizations face when designing and implementing documentation practices. These considerations include balancing customization with standardization, deciding between single or multiple forms of documentation, determining the extent of detail to include, choosing the appropriate level of automation to employ in the documentation process, and selecting between static or interactive formats. Based on our analysis, we offer recommendations to help organizations develop documentation strategies that align with their governance goals and stakeholder needs.
The primary contribution of our research is that our synthesis identifies patterns, gaps, and emerging trends across individual publications about documentation approaches and evaluations. By offering a holistic analysis of how AI documentation can support robust governance in real-world contexts, we clarify which documentation practices are likely to be most effective, under what conditions, and for which stakeholders. As such, our research can serves as a valuable resource for both researchers and practitioners seeking to refine and enhance their approaches to AI documentation in applied AI development contexts.
@@ 2. What organizations could document about AI systems
Model documentation also enhances AI development and governance, providing a window into a model’s creation, performance, and risks (Crisan et al., 2022; McMillan-Major et al., 2021; Mitchell et al., 2019; Shen et al., 2022; Stoyanovich and Howe, 2019). Documenting aspects such as the model’s architecture, training procedures, intended use cases, and evaluation methods enables practitioners to assess potential risks and responsibly adapt models to specific contexts. For models developed in stages—such as those pretrained on large datasets and later fine-tuned—documenting each stage is essential. This helps downstream users understand the development pipeline and its implications for their intended use.
AI services or applications in the real world typically consist of complex systems rather than isolated models. These systems often include multiple components, such as machine learning models, trust and safety filters, third-party software integrations, and user interfaces. The interaction among these components significantly affects the system’s overall performance and risk profile — especially because the properties or behavior of individual components may change when integrated into a system (Smart et al., 2024). Recognizing that AI systems are more than just the sum of their parts, researchers have suggested that organizations should document not only each individual component but also the system as a whole (Arnold et al., 2019; Baracaldo et al., 2022; Brajovic et al., 2023; Green et al., 2021; Procope et al., 2022; Yang et al., 2018; Blasch et al., 2021). System documentation could provide an overview, including objectives, inputs, outputs, and a diagram illustrating how the components interact. This comprehensive approach ensures that practitioners have a clear understanding of the system’s behavior in its entirety.
@@
We can categorize documentation stakeholders into two broad groups: documentation producers and documentation consumers. Documentation producers are those who actively contribute to the generation of documentation artifacts, while documentation consumers are those who read and use it. While this grouping simplifies some important distinctions, it serves as a helpful heuristic for understanding the opportunities and tensions between groups with both shared and distinct goals. We consider both producers and consumers and their interplay in our analysis.
We also note that ”AI documentation” could refer to either the process of producing documentation or the resulting artifacts. While documentation artifacts inform stakeholders about the responsible use of AI, the process of creating these documents can also institutionalize best practices and foster a culture of risk management.. Therefore, when considering how documentation can improve governance outcomes, we address both artifacts and processes.
@@ 3.1. Publication Sampling
We employed a purposeful sampling approach (Palinkas et al., 2015) to identify proposed documentation frameworks and relevant empirical studies. Given our specific interest in how AI documentation can support governance, our use of a purposeful sampling approach allowed us to focus on proposed frameworks and empirical studies that are directly relevant to our research questions. It also afforded us the flexibility to iteratively refine our sample of papers based on emerging findings.
@@
We identified an initial seed sample of publications from a systematic review focused on AI documentation methods relevant to EU regulation (Micheli et al., 2023). We cross-referenced this initial sample against the references used by the Partnership on AI in developing their ABOUT ML framework,333https://partnershiponai.org/workstream/about-ml/ which seeks to provide technology organizations with guidance on what to document about their AI systems. We chose these initial sources because they allowed us to focus on publications that aimed to establish best practices for industry documentation and address policy initiatives geared toward institutionalizing these practices. We reviewed the cited references within each of these initial works for additional proposed documentation frameworks and empirical studies related to documentation. We also consulted with academic researchers and industry practitioners with experience producing and using documentation to identify additional frameworks and studies. We excluded papers that centered on tools for implementing documentation (e.g., automated documentation code libraries) or that were not specifically focused on AI system documentation, such as fairness checklists.
@@
This process yielded a sample of 37 proposed frameworks, eight of which included empirical evidence related to their implementation. Because of the relative dearth of empirical findings within publications proposing approaches, we further expanded our sample by searching for references within these 37 approaches that empirically evaluated documentation methods. We also conducted searches in the archives of the ACM Conference on Human Factors in Computing Systems and the ACM Conference on Computer Supported Cooperative Work using terms such as ”datasheet,” ”model card,” ”AI documentation,” ”model documentation,” ”data card,” and ”data documentation.” Together, these methods allowed us to identify 13 additional publications with empirical findings.
Recognizing the importance of applied perspectives, we included non-peer-reviewed papers, such as technical whitepapers published by industry actors, acknowledging that they often provide insights not found in academic literature. While our sampling approach was not exhaustive, we reached theoretical saturation after analyzing the selected publications, as no new themes emerged. Therefore, we concluded our sampling at this point.
@@ 3.2. Data Analysis
Our data analysis followed an abductive approach (Timmermans and Tavory, 2012; Tavory and Timmermans, 2014), iterating between inductively derived codes, as in grounded theory methods (Strauss and Corbin, 1990; Charmaz, 2014), and codes based on the theoretical motivations of our research. The first author conducted the initial coding by applying descriptive codes aligned with our research goals, such as identifying how AI documentation enhances internal governance and the challenges of using documentation as a governance tool. After discussing the results with the second author, we refined these codes into higher-order categories during the axial coding phase. For instance, codes like ”identify intended purpose” and ”specify out-of-scope uses” were grouped under the category ”inform downstream documentation consumers.”
In the final thematic coding phase, we grouped axial codes into broader theoretical mechanisms through which AI documentation could support robust governance, the challenges in achieving these impacts, and the design trade-offs faced by organizations.
@@ 4.1. Publication Sample
We also categorized the type of evaluation each framework or independent empirical research study employed. Frameworks employing a ”feasibility analysis” are those where the framework’s authors or another group applied the framework to create documentation for a hypothetical or actual dataset, model, system, or method. This type of analysis demonstrates that the framework could theoretically be used for its intended purpose but does not involve empirical evaluation with practitioners in research or real-world settings. If a study developed a documentation artifact for the purpose of an empirical study, we classified this as part of the empirical study rather than as a feasibility analysis, as empirical studies offer a more rigorous evaluation.
@@
In some instances, framework authors mentioned consulting relevant stakeholders during the design or refinement of their framework. However, if these consultations are only briefly mentioned or not elaborated upon, we do not classify the work as including a practitioner study.
@@
| 2 | Arnold et al. | 2019 | System | Feasibility analysis |
| 3 | Baracaldo et al. | 2022 | System | Feasibility analysis |
| 4 | Bender & Friedman | 2018 | Data | Feasibility analysis |
| 5 | Bhat et al. | 2023 | Evaluation only | Practitioner lab study |
| 6 | Blasch et al. | 2020 | System | Feasibility analysis |
| 7 | Boyd | 2021 | Evaluation only | Practitioner lab study |
| 8 | Brajovic et al. | 2023 | System | Feasibility analysis |
| 9 | Chang & Custis | 2022 | Evaluation only | Practitioner real-world study |
| 10 | Chmielinski et al. | 2020 | Data | Feasibility analysis |
| 11 | Chmielinski et al. | 2024 | Method, process, or task | None |
| 12 | Crisan et al. | 2022 | Model | Feasibility analysis, Practitioner lab study |
| 13 | Díaz et al. | 2023 | Data | Feasibility analysis |
| 14 | Gebru et al. | 2021 | Data | Feasibility analysis |
| 15 | Geiger et al. | 2020 | Evaluation only | Artifact study |
| 16 | Gilbert et al. | 2023 | System | Feasibility analysis |
@@
| 17 | Heger et al. | 2022 | Evaluation only | Practitioner real-world study |
@@
| 18 | Hind et al. | 2019 | Evaluation only | Practitioner real-world study |
| 19 | Holland et al. | 2018 | Data | Feasibility analysis |
| 20 | Hupont & Gomez | 2022 | Method, process, or task | Feasibility analysis |
| 21 | Liang et al. | 2024 | Evaluation only | Artifact study |
| 22 | Liao et al. | 2023 | Evaluation only | Practitioner lab study |
| 23 | Marone & Van Durme | 2023 | Data | Feasibility analysis |
| 24 | McMillan-Major et al. | 2021 | Data, Model | Feasibility analysis |
| 25 | McMillan-Major, Bender, & Friedman | 2024 | Data | Practitioner lab study |
| 26 | Miceli et al. | 2021 | Evaluation only | Practitioner real-world study |
| 27 | Mitchell et al. | 2019 | Model | Feasibility analysis |
| 28 | Mohammad | 2022 | Method, process, or task | Feasibility analysis |
| 29 | Moore, LIao, & Subramonyam | 2023 | Evaluation only | Practitioner lab study |
| 30 | Nunes et al. | 2022 | Evaluation only | Practitioner lab study |
| 31 | Papakyriakopoulos et al. | 2023 | Data | Feasibility analysis |
| 32 | Pepe et al. | 2024 | Evaluation only | Artifact study |
| 33 | Procope et al. | 2022 | System | Feasibility analysis |
| 34 | Pushkarna et al. | 2022 | Data | Practitioner real-world study |
| 35 | Raji & Yang | 2020 | System | None |
@@
| 36 | Reid & Williams | 2023 | Evaluation only | Practitioner real-world study, Artifact study |
| 37 | Richards et al. | 2020 | System | Feasibility analysis |
| 38 | Roman et al. | 2023 | Data | Feasibility analysis |
| 39 | Rostamzadeh et al. | 2022 | Data | Practitioner lab study |
| 40 | Shen et al. | 2022 | Model | Practitioner lab study |
| 41 | Shimorina & Belz | 2021 | Method, process, or task | None |
| 42 | Soh | 2021 | Data | None |
| 43 | Sokol & Flach | 2020 | Method, process, or task | Feasibility analysis |
| 44 | Srinivasan et al. | 2021 | Data | Feasibility analysis |
| 45 | Stoyanovich & Howe | 2019 | Data, Model | Feasibility analysis |
| 46 | Subramaniam et al. | 2023 | Data | Practitioner lab study |
| 47 | Sun et al. | 2019 | Data | Feasibility analysis |
| 48 | Tagliabue et al. | 2021 | Method, process, or task | Feasibility analysis |
| 49 | Yang et al. | 2018 | Method, process, or task | Feasibility analysis |
| 50 | Zheng et al., | 2022 | Data | Practitioner lab study |
4.2. Potential Impacts & Challenges
While empirical evidence directly linking documentation artifacts or processes to improved governance outcomes is limited, the theories proposed in existing documentation frameworks offer compelling hypotheses about how documentation can positively impact governance. Empirical studies involving practitioners, alongside analyses of publicly available documentation artifacts, provide an initial glimpse into the factors that contribute to accurate, high-quality, and comprehensive documentation, as well as the reasons why documentation efforts may sometimes fall short of their goals.
Our analysis identified four possible ways in which documentation could improve governance: informing downstream users about system development and associated risks, encouraging ethical reflection among practitioners, facilitating communication among stakeholders, and enhancing AI development and governance overall. In the following sections, we explore each of these hypothesized impacts in detail and examine the challenges real-world organizations face in realizing these benefits.

### v2

@@ Abstract.
Documentation plays a crucial role in both external accountability and internal governance of AI systems. Although there are many proposals for documenting AI data, models, systems, and methods, the ways these practices enhance governance as well as the challenges practitioners and organizations face with documentation remain underexplored. In this paper, we analyze 37 proposed documentation frameworks and 22 empirical studies evaluating their use. We identify several pathways or ”theories of change” through which documentation can enhance governance, including informing stakeholders about AI risks and applications, facilitating collaboration, encouraging ethical deliberation, and supporting best practices. However, empirical findings reveal significant challenges for practitioners, such as insufficient incentives and resources, structural and organizational communication barriers, interpersonal and organizational constraints to ethical action, and poor integration with existing workflows. These challenges often hinder the realization of the possible benefits of documentation. We also highlight key considerations for organizations when designing documentation, such as determining the appropriate level of detail and balancing automation in the process. We conclude by discussing how future research can expand on our findings such as by exploring documentation approaches that support governance of general-purpose models and how multiple transparency and documentation methods can collectively improve governance outcomes.
@@ 1. Introduction
Transparency in AI is widely regarded as essential for calibrating trust and supporting accountability (Liao and Vaughan, 2024). By providing stakeholders with a clear view of an AI system’s composition, operation, and development process, transparency allows for informed oversight and critical evaluation. When AI systems are thoroughly documented, these documentation artifacts offer invaluable insights into important system features such as training data, algorithms, and risk management strategies. This information can support several groups of external stakeholders. Documentation can help downstream deployers and users understand how the system functions and what risks it conveys. It can also assist policymakers and researchers in holding companies accountable for the negative consequences of their AI technologies.
Yet the importance of documentation extends beyond external accountability.111Though the concepts of documentation and transparency are often blurred, complete documentation about an AI system does not guarantee that information will be made available to interested stakeholders, nor do transparency artifacts always reflect the full or most relevant details of an AI system even though such details may have been captured in internal documents. We distinguish between the concepts in this report in order to provide greater clarity about the presumptive goals of documentation and an exploration of what practices may more likely support those objectives. Internally, documentation serves as a critical tool for managing AI systems throughout their lifecycle for a wide range of stakeholders. For example, documentation could help downstream technical practitioners understand the strengths, limitations, and risks of training a model on a given dataset. It might also help compliance and governance teams assess a potential system use case’s compliance with company policies or legal requirements. Or the process of producing documentation might lead practitioners to make different decisions about how to sufficiently mitigate potential harms. Robust documentation can improve the efficiency and quality of other important AI governance practices and processes, such as internal and third-party auditing, harm identification, risk triaging, and compliance functions, and can help practitioners better implement risk management practices, such as system evaluation and risk mitigation approaches. In this way, documentation serves as a foundational building block of a multifaceted AI governance approach, both providing important inputs to other governance processes and creating a scaffold on which additional governance practices can build over time.
@@
In this paper, we explore how documentation processes and artifacts can serve as one critical building block in supporting AI governance and risk management goals. Our research draws from a qualitative analysis of academic and gray literature,222Gray literature refers to publications that have not been peer-reviewed, but that present detailed theories or research findings including 37 proposed documentation frameworks and 22 papers with relevant empirical findings. From this analysis, we identify four ”theories of change” or pathways through which documentation could improve AI governance, including informing stakeholders about responsible use, facilitating collaboration on AI risks, encouraging ethical deliberation, and improving overall governance and development practices. Our analysis also identified several challenges that might prevent documentation from effectively supporting governance aims. These include insufficient incentivization and resourcing for documentation, structural and organizational barriers to effective collaboration, interpersonal and organizational constraints to ethical action, and poor integration with practitioners’ day-to-day workflows and processes.
We also explored considerations or trade-offs organizations face when designing and implementing documentation practices to balance competing priorities or documentation goals. These considerations include balancing customization with standardization, deciding between single or multiple forms of documentation to address different audiences, determining the extent of detail to include, choosing the appropriate level of automation to employ in the documentation process, and selecting between static or interactive formats.
The primary contribution of our research is that our synthesis identifies patterns, gaps, and emerging trends across individual publications about documentation approaches and evaluations. By offering a holistic analysis of how AI documentation can support robust governance in real-world contexts, we clarify which documentation practices are likely to be most effective, under what conditions, and for which stakeholders. As such, our research can serves as a valuable resource for both researchers and practitioners seeking to refine and enhance their approaches to AI documentation in applied AI development contexts. We conclude with a discussion of how future research can build on or expand upon the themes we explore throughout the paper.
@@ 2. What organizations could document about AI systems
Model documentation also enhances AI development and governance, providing a window into a model’s creation, performance, and risks (Crisan et al., 2022; McMillan-Major et al., 2021; Mitchell et al., 2019; Shen et al., 2022; Stoyanovich and Howe, 2019). Documenting aspects such as the model’s architecture, training procedures, intended use cases, and evaluation methods enables practitioners to assess potential risks and responsibly adapt models to specific contexts. For models developed in stages—such as those pretrained on large datasets and later fine-tuned—documenting each stage is essential (Gilbert et al., 2023). This helps downstream users understand the development pipeline and its implications for their intended use.
AI services or applications in the real world typically consist of complex systems rather than isolated models. These systems often include multiple components, such as machine learning models, trust and safety filters, third-party software integrations, and user interfaces. The interaction among these components significantly affects the system’s overall performance and risk profile — especially because the properties or behavior of individual components may change when integrated into a system (Smart et al., 2024). Recognizing that AI systems are more than just the sum of their parts, researchers have suggested that organizations should document not only each individual component but also the system as a whole (Arnold et al., 2019; Baracaldo et al., 2022; Brajovic et al., 2023; Procope et al., 2022; Yang et al., 2018; Blasch et al., 2021). System documentation could provide an overview, including objectives, inputs, outputs, and a diagram illustrating how the components interact (Procope et al., 2022). This comprehensive approach ensures that practitioners have a clear understanding of the system’s behavior in its entirety.
@@
We can categorize documentation stakeholders into two broad groups: documentation producers and documentation consumers. Documentation producers are those who actively contribute to the generation of documentation artifacts, while documentation consumers are those who read and use it. This grouping simplifies some important distinctions.333Practitioners sometimes create documentation for their own use, while in other instances, they produce one type of documentation for others and rely on a different type created by someone else. For example, a compliance team might use documentation from a technical team to understand how a system was developed and functions in order to assess compliance concerns. At the same time, the compliance team might generate documentation for technical teams, outlining the limitations that apply to the use of a system or component. In this way, the same stakeholder could be both a consumer and a producer of documentation, even for the same system or component. Still, distinguishing between consumers and producers serves as a helpful heuristic for understanding the opportunities and tensions between groups with both shared and distinct goals. We consider both producers and consumers and their interplay in our analysis.
We also note that ”AI documentation” could refer to either the process of producing documentation or the resulting artifacts. While documentation artifacts can inform stakeholders about the responsible use of AI, the process of creating these documents can also institutionalize best practices and foster a culture of risk management. Therefore, when considering how documentation can improve governance outcomes, we address both artifacts and processes.
@@ 3.1. Publication Sampling
We employed a purposeful sampling approach (Palinkas et al., 2015; Suri, 2011) to identify publications. This method was particularly well-suited to our study because it allowed us to target information-rich frameworks and empirical studies clearly aligned with our research goals of understanding the functions documentation plays in supporting governance and the challenges associated with using documentation as a foundational governance tool. By prioritizing publications that offered insights into documentation frameworks, evaluation methods, and practical applications, we ensured extensive coverage of existing theories and practices while maintaining direct relevance to our research questions. Furthermore, our iterative sampling process enabled us to refine our dataset as our analysis progressed, incorporating studies that contributed meaningful data or theoretical perspectives. Our approach contrasts with systematic sampling, which establishes strict inclusion and exclusion criteria beforehand. While systematic sampling might have offered more exhaustive coverage, it likely would have included less relevant papers that did not align as closely with our objectives and would not have allowed us to expand our sample in response to emerging findings.
We opted to include both proposed frameworks and empirical evaluations in our analysis rather than limiting our focus to empirical studies. This decision stemmed from the recognition that the goals of a documentation framework and its real-world implementation can often diverge. For instance, a framework that generates high-quality documentation in a research context may fall short when applied in practical settings. By examining both the intended goals of proposed frameworks and the empirical evidence, we aimed to better understand and highlight the gap between these two dimensions, informing future designs and evaluations.
@@
We identified an initial seed sample of publications from a systematic review focused on AI documentation methods relevant to EU regulation (Micheli et al., 2023). We cross-referenced this initial sample against the references used by the Partnership on AI in developing their ABOUT ML framework,444https://partnershiponai.org/workstream/about-ml/ which seeks to provide technology organizations with guidance on what to document about their AI systems. We chose these initial sources because they allowed us to focus on publications that aimed to establish best practices for industry documentation and address policy initiatives geared toward institutionalizing these practices. We reviewed the references of these publications and consulted with subject matter experts in industry and academia with experience producing or using documentation to identify additional publications.
@@
This process yielded a sample of 37 proposed frameworks, eight of which included empirical evidence related to their implementation. To expand our sample of empirical findings, we again consulted with subject matter experts and reviewed the citations of the frameworks for references to further empirical work. We also conducted searches in the archives of the ACM Conference on Human Factors in Computing Systems and the ACM Conference on Computer Supported Cooperative Work using terms such as ”datasheet,” ”model card,” ”AI documentation,” ”model documentation,” ”data card,” and ”data documentation.” Together, these methods allowed us to identify 14 additional publications with empirical findings. Our sampling concluded upon reaching theoretical saturation, where further sampling no longer yielded new themes, supporting the adequacy and richness of our chosen publications.
Recognizing the importance of applied perspectives, we included non-peer-reviewed papers, such as technical whitepapers published by industry actors, acknowledging that they often provide insights not found in academic literature. To maintain a targeted focus on publications examining the governance role of documentation rather than the technical aspects of specific tools, we excluded papers discussing enabling technologies (e.g., automated documentation libraries or supporting software) or those addressing broader topics like fairness checklists not specific to AI documentation.
@@ 3.2. Data Analysis
We employed an abductive approach for our data analysis (Tavory and Timmermans, 2014; Timmermans and Tavory, 2012), a qualitative method that involves alternating between theoretical concepts identified deductively and codes generated inductively through data review, similar to grounded theory methods. (Charmaz, 2014; Strauss and Corbin, 1990; Glaser, 1978; Glaser et al., 1968). Guided by this approach, we initially applied codes aligned with our objectives of understanding how documentation can support robust AI governance and identifying barriers to its effectiveness. At the same time, we remained open to adjusting our coding focus based on insights gained through inductive analysis of our publication sample. Our analysis was refined iteratively through ongoing discussions and repeated reviews of the sample publications in alignment with these objectives.
In the initial coding phase, the first author applied descriptive codes to papers that aligned with our research objectives. These codes included terms like “discoverability,” “standardization,” “intended use,” “incentivization,” “documentation errors.” During the review of these descriptive codes, both authors observed emerging themes related to both theories of change and challenges. For example, we noted that both the codes ”identify intended purpose” and ”specify out of scope uses” both pertained to informing downstream documentation consumers. Following this discussion, we organized our descriptive codes into four theory of change categories: informing downstream stakeholders, facilitating collaboration and communication, promoting ethical deliberation, and supporting governance and development best practices. We grouped codes related to challenges into four overarching categories: insufficient incentives and resources, structural and organizational obstacles to effective communication, limited ethical awareness and barriers to ethical action, and inadequate integration with existing workflows and organizational practices. We note that in some cases, challenges are more relevant to specific theories of change, and in other cases, challenges are more generally applicable. We discuss these points of connection in the results.
The inductive aspects of our analysis highlighted that authors of documentation publications made varied design choices aligned with the goals of their proposed frameworks. Consequently, we expanded our focus to examine the design tensions and considerations that surfaced within and across these papers. To capture these dynamics, we grouped descriptive codes related to design choices, particularly those that appeared to exist in tension with one another. These considerations included the balance between customization and standardization in addressing specific components or use cases, the extent of audience-specific tailoring in documentation artifacts, the level of interactivity provided, the amount of detail included, and the degree of automation in the documentation process.
3.3. Positionality
The authors of this study work at a civil society organization specializing in technology policy. Previously, the first author developed machine learning models, and the second author focused on AI governance and policy within technology companies. Both authors have experience conducting research in human-centered and responsible AI and have advised government and industry on implementing risk management practices for AI systems. Our academic backgrounds in psychology and public policy inform our understanding of AI as a sociotechnical system and AI governance as a sociotechnical process. These perspectives inevitably shaped our selection of publications, the themes we chose to explore, and our synthesis of the results. Given that our analysis approach is interpretive, researchers with different perspectives and backgrounds may have arrived at different conclusions when analyzing the same dataset or found different publications more meaningful to the overall research aims.
@@ 4.1. Publication Sample
We also categorized the type of evaluation each framework or empirical research study employed. Frameworks employing a ”feasibility analysis” are those where the framework’s authors or another group applied the framework to create documentation for a hypothetical or actual dataset, model, system, or method. This type of analysis demonstrates that the framework could theoretically be used for its intended purpose but does not involve empirical evaluation with practitioners in research or real-world settings. If a study developed a documentation artifact for the purpose of an empirical study, we classified this as part of the empirical study rather than as a feasibility analysis, as empirical studies offer a more rigorous evaluation.
@@
In some instances, framework authors mentioned consulting relevant stakeholders during the design or refinement of their framework. However, if these consultations were only briefly mentioned, we do not classify the work as including a practitioner study.
@@
| 2 | Ahlawat, Winecoff, & Mayer | 2024 | Evaluation only | Practitioner real-world study |
| 3 | Arnold et al. | 2019 | System | Feasibility analysis |
| 4 | Baracaldo et al. | 2022 | System | Feasibility analysis |
| 5 | Bender & Friedman | 2018 | Data | Feasibility analysis |
| 6 | Bhat et al. | 2023 | Evaluation only | Practitioner lab study |
| 7 | Blasch et al. | 2020 | System | Feasibility analysis |
| 8 | Boyd | 2021 | Evaluation only | Practitioner lab study |
| 9 | Brajovic et al. | 2023 | System | Feasibility analysis |
@@
| 10 | Chang & Custis | 2022 | Evaluation only | Practitioner real-world study |
| 11 | Chmielinski et al. | 2020 | Data | Feasibility analysis |
| 12 | Chmielinski et al. | 2024 | Method, process, or task | None |
| 13 | Crisan et al. | 2022 | Model | Practitioner lab study |
| 14 | Díaz et al. | 2023 | Data | Feasibility analysis |
| 15 | Gebru et al. | 2021 | Data | Feasibility analysis |
| 16 | Geiger et al. | 2020 | Evaluation only | Artifact study |
| 17 | Gilbert et al. | 2023 | System | Feasibility analysis |
| 18 | Heger et al. | 2022 | Evaluation only | Practitioner real-world study |
| 19 | Hind et al. | 2019 | Evaluation only | Practitioner real-world study |
| 20 | Holland et al. | 2018 | Data | Feasibility analysis |
| 21 | Hupont & Gomez | 2022 | Method, process, or task | Feasibility analysis |
| 22 | Liang et al. | 2024 | Evaluation only | Artifact study |
| 23 | Liao et al. | 2023 | Evaluation only | Practitioner lab study |
| 24 | Marone & Van Durme | 2023 | Data | Feasibility analysis |
| 25 | McMillan-Major et al. | 2021 | Data, Model | Feasibility analysis |
| 26 | McMillan-Major, Bender, & Friedman | 2024 | Data | Practitioner lab study |
| 27 | Miceli et al. | 2021 | Evaluation only | Practitioner real-world study |
| 28 | Mitchell et al. | 2019 | Model | Feasibility analysis |
| 29 | Mohammad | 2022 | Method, process, or task | Feasibility analysis |
| 30 | Moore, LIao, & Subramonyam | 2023 | Evaluation only | Practitioner lab study |
| 31 | Nunes et al. | 2022 | Evaluation only | Practitioner lab study |
| 32 | Papakyriakopoulos et al. | 2023 | Data | Feasibility analysis |
| 33 | Pepe et al. | 2024 | Evaluation only | Artifact study |
| 34 | Procope et al. | 2022 | System | Feasibility analysis |
| 35 | Pushkarna et al. | 2022 | Data | Practitioner real-world study |
| 36 | Raji & Yang | 2020 | System | None |
@@
| 37 | Reid & Williams | 2023 | Evaluation only | Practitioner real-world study, Artifact study |
| 38 | Richards et al. | 2020 | System | Feasibility analysis |
| 39 | Roman et al. | 2023 | Data | Feasibility analysis |
| 40 | Rostamzadeh et al. | 2022 | Data | Practitioner lab study |
| 41 | Shen et al. | 2022 | Model | Practitioner lab study |
| 42 | Shimorina & Belz | 2021 | Method, process, or task | None |
| 43 | Soh | 2021 | Data | None |
| 44 | Sokol & Flach | 2020 | Method, process, or task | Feasibility analysis |
| 45 | Srinivasan et al. | 2021 | Data | Feasibility analysis |
| 46 | Stoyanovich & Howe | 2019 | Data, Model | Feasibility analysis |
| 47 | Subramaniam et al. | 2023 | Data | Practitioner lab study |
| 48 | Sun et al. | 2019 | Data | Feasibility analysis |
| 49 | Tagliabue et al. | 2021 | Method, process, or task | Feasibility analysis |
| 50 | Yang et al. | 2018 | Method, process, or task | Feasibility analysis |
| 51 | Zheng et al., | 2022 | Data | Practitioner lab study |
4.2. Theories of change
While empirical evidence directly linking documentation artifacts or processes to improved governance outcomes is limited, proposed documentation frameworks offer compelling theories of change about how documentation can positively a multifaceted impact governance approach. Our analysis revealed four such theories of change: informing downstream users about system development and associated risks, encouraging ethical reflection among practitioners, facilitating communication among stakeholders, and enhancing AI development and governance overall.